OpenStudy (theraggedydoctor):

If \(\frac{ 1 }{ 2 }\) is \(\frac{ 3 }{ 4 }\) of \(\frac{ 4 }{ 5 }\) of a certain number, what is that number?

OpenStudy (michele_laino):

I think that your problem can be modeled by this equation: \[\Large \frac{3}{4} \times \frac{4}{5} \times N = \frac{1}{2}\] where N is the requested number